2. A block of 3 kg is sitting at rest on an incline of 10 degrees. A constant pushing force of 40 N is applied in the uphill direction for 1.3 seconds. The incline also provides a constant friction force of 20 N. Use the impulse-momentum approach to calculate how long the block will continue to slide before it comes to a stop on the incline.
Answer #1

Net force - F. F-40N may sina ingesa mg along the inclined for fost 1.3 second- Enet E-mgsino-fr = 40-3x9.8 xSino) -20 = 14.8I m ( 22 - 262) =) z 2 2 501) - Fnet x8 (-ve sign es used in Lats because not force és opposite tie duflacement) -25.1188 =